OpenStudy (anonymous):

3.Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing.

OpenStudy (anonymous):

Here is what I have for the other 2 questions: 1. Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences. F(x) = 15(.3)^x This equation would help you find out the snails population each year. 2. A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population. G(x) = 200(1.03)^2x 1.03 x 1.03 = 1.0609 G(x) = 200(1.0609)^x